# Quantum memory assisted observable estimation

### v3

@@ Abstract
The estimation of many-qubit observables is an essential task of quantum information processing. The generally applicable approach is to decompose the observables into weighted sums of multi-qubit Pauli strings, i.e., tensor products of single-qubit Pauli matrices, which can readily be measured with single qubit rotations. The accumulation of shot noise in this approach, however, severely limits the achievable variance for a finite number of measurements. We introduce a novel method, dubbed coherent Pauli summation (CPS), that circumvents this limitation by exploiting access to a single-qubit quantum memory in which measurement information can be stored and accumulated. Our algorithm offers a reduction in the required number of measurements for a given variance that scales linearly with the number of Pauli strings in the decomposed observable. Our work demonstrates how a single long-coherence qubit memory can assist the operation of many-qubit quantum devices in a cardinal task.
@@ 1 Introduction
One approach for observable estimation with minimal sampling complexity is the quantum phase estimation (QPE) algorithm [16, 15, 37, 26, 41, 29]. Its implementation, however, requires qubit systems with low noise and long coherence times for high precision estimation since the measurement circuit depth is inversely proportional to the achievable precision.
As an alternative, the quantum energy (expectation) estimation (QEE) approach of decomposing the observable into a weighted sum of $N$ multi-qubit Pauli strings is commonly used in the variational quantum eigensolver [31]. While QEE minimizes the measurement circuit depth by requiring only a single layer of single-qubit rotations, it also suffers from increased sample complexity due to the accumulation of shot noise in the estimation procedure. Specifically, the expectation values of Pauli strings are estimated independently, and the observable is then calculated as a linear combination of these.
@@
Consequently, to estimate an observable comprised of $N$ Pauli stings to a variance $\eta$, each Pauli string should be estimated to a variance $O(\eta/N)$ resulting in an overall sample complexity scaling as $O(N^{2})$. This accumulation of noise poses a “shot noise bottleneck" since the amount of measurements will ultimately be limited by the available runtime of the device before, e.g., re-calibration of the device is needed. The measurement process itself is also often one of the most time consuming operations in current quantum devices [14, 4, 35].
@@
Here, we propose a novel approach dubbed coherent Pauli summation (CPS) that overcomes the shot-noise bottleneck through the use of a single-qubit quantum memory (QM). Our method allows for a direct measurement of the multi-qubit observable by estimating the phase of the memory qubit at the end of the protocol. Hence, the accumulation of shot noise, originating from the summation of individually estimated mean values of Pauli strings in the QEE approach, is prevented. Importantly, this is obtained with a measurement circuit depth that scales only logarithmically with the required estimation variance, in contrast to the linear scaling of the QPE algorithm.
The critical point of CPS is to employ quantum signal processing (QSP) techniques to encode the mean value of Pauli strings in the phase of a single qubit quantum memory. The lack of shot noise accumulation results in a gain in the variance of the estimate compared to the QEE of $O(1/N)$, where $N$ is the number of Pauli strings in the observable decomposition. This gain can be understood as obtaining a Heisenberg limited scaling of the variance rather than the standard quantum limit scaling of the QEE. The necessary coherence time of the single memory qubit required for CPS scales linearly with $N$, while the required coherence time overhead of the quantum device only increases logarithmically with the required variance. The CPS thus outlines how a single long-coherence qubit memory can advance the performance of larger scale quantum devices.
@@ 2 CPS method
In general, the state $|\Psi\rangle$ is repeatedly prepared, and the Pauli strings are measured sequentially via projective measurements to obtain estimates of every Pauli string: $\langle\hat{P}_{j}\rangle\equiv\langle\Psi|\hat{P}_{j}|\Psi\rangle$ [31, 24]. The mean value of the observable $\langle\hat{O}\rangle$ can be calculated classically after estimating all $\langle\hat{P}_{j}\rangle$, $j=1\ldots N$. This approach, however, suffers from the accumulation of shot noise from the individually estimated mean values of the Pauli strings, as described above. If each estimate $\langle\hat{P}_{j}\rangle$ is estimated with a variance $\sigma^{2}(\langle\hat{P}_{j}\rangle)\sim\eta_{p}$, the variance of the final estimate is $\sigma^{2}(\hat{\langle O\rangle})\sim N\eta_{p}$, assuming roughly equal weights of the $\hat{P}_{j}$ in the decomposition of $\hat{O}$.
@@
    Obtain rough estimates of the mean values of every Pauli string $\langle\Psi_{0}|\hat{P}_{j}|\Psi_{0}\rangle=\langle\hat{P}_{j}\rangle$, $j=1,\dots,N$ by performing $O\left(\log{\left(\frac{N}{\sqrt{\eta}}\right)}\right)$ projective measurements similar to the QEE approach. This step estimates $s_{j}=\text{sign}{(a_{j}\langle\hat{P}_{j}\rangle)}$ i.e. the sign information of the Pauli strings.
@@
    In the second step, $s_{j}\epsilon|a_{j}\langle\hat{P}_{j}\rangle|$ is directly encoded in the phase of the single qubit memory using a modified phase-kickback algorithm [18] together with QSP techniques. Here, $\epsilon\in(0,1)$ is a tuning parameter that is used to circumvent the general "modulo $2\pi$" ambiguity of phase estimation, which we will detail below. The encoding is done sequentially for all Pauli strings in the decomposition, resulting in a final phase of $\sim\epsilon\langle\hat{O}\rangle$ followed by a projective measurement of the single qubit state.
@@
    The previous step is repeated a number of times with a varying parameter $\epsilon$ to obtain the final estimate of $\langle\hat{O}\rangle$. In particular, the procedure is repeated $M_{l}=\alpha+\gamma(d_{L}-l)$ times with $\epsilon\propto 2^{l}$ for $l=1,2,\dots,d_{L}$, $d_{L}=[\log_{2}{1/\eta}]$. The increasing powers of $\epsilon$ enable estimation of the digits of $\langle\hat{O}\rangle$, thereby circumventing the $2\pi$ ambiguity of the phase estimation. Analysing the upper bound of the variance of the estimate, in [15], the Heisenberg scaling with the best constant can be achieved by selecting $\alpha=3$ and $\gamma=1$ without any extra assumptions on $M_{l}$. In general, the Heisenberg scaling is obtained for any $\alpha>2$ and $\gamma>0$. For more details, please see Appendix B. The total amount of repetitions of the second step with different powers of $\epsilon$ is $M_{q}=\sum_{l=0}^{d_{L}}M_{l}$.
At the end of these three steps, $\langle\Psi_{0}|\hat{O}|\Psi_{0}\rangle$ can be estimated up to a variance of $\eta$ with an overall sample complexity of $N/\eta$. We have counted this scaling as the number of state preparation circuits $V$ required in the method, which also includes the extra state preparation circuits that are part of the QSP step as detailed below.
@@
Figure 2: Quantum circuit realising the QSP by a sequence of controlled unitaries $\hat{U}_{\phi_{i}}(\theta_{j})$, $i=1,\dots,n$, built from Hadamard gates, $\hat{R}_{\phi}=e^{\frac{i\phi\hat{\sigma}_{z}}{2}}$, and controlled $\hat{U}_{P_{j}}$, |$\ket{\psi}_{m}=\alpha\ket{+}_{m}+\beta\ket{-}_{m}$ denotes the state of a single memory qubit.
Method Number of state preparations Qubits Coherence times
QEE $N^{2}/\eta$ Processing qubits $t_{prep}$
CPS $(N/\eta)O\left(\frac{\log{(N/\sqrt{\eta})}}{\log{\log{(N/\sqrt{\eta})}}}\right)$ Processing qubits $t_{prep}O(\log{(N/\sqrt{\eta})})$
Memory qubit $Nt_{prep}O\left(\frac{\log{(N/\sqrt{\eta})}}{\log{\log{(N/\sqrt{\eta})}}}\right)$
Table 1: Comparison of resources for the QEE and CPS methods for a fixed target variance $\eta$. We have quantified the necessary coherence times of the processing qubits of quantum device and the memory qubit in terms of the state preparation time, $t_{prep}$, which is assumed to be the dominant timescale. The observable is assumed to be decomposed into a summation of $N$ Pauli strings.
@@ 2.1 Phase wrapping
Since the phase encoding only provides an estimate of $\langle\hat{O}\rangle$ mod $2\pi$, we are facing a phase wrapping problem. To mitigate this issue, we have included the factor of $\epsilon$ in the encoding. This brings us to the third and final step of the CPS. To control the phase wrapping, we use the sampling approach introduced in Refs. [10, 15]. Instead of using fixed $\epsilon$ to encode $\theta_{j}$, we sample at multiple orders $\tilde{\epsilon}_{l}\equiv 2^{l}\epsilon$, $l=1,2,\dots,d_{L}$ to gradually enclose on $\langle\hat{O}\rangle$. This trick can be seen as sequentially estimating the digits of $\langle\hat{O}\rangle$. The parameter $\epsilon$ is selected in a way that $\sum_{j=1}^{N}s_{j}\epsilon|a_{j}|\cos{\theta_{j}}<2\pi$, holds. By a proper selection of $d_{L}$ and the number of measurements, we find that the CPS described above results in an estimate of $\hat{O}$ with a variance of $\sigma^{2}(\hat{O}_{CPS})\sim NT^{-1}$, where $T$ is the total number of state preparation circuits ($\hat{V}$ and $\hat{V}^{\dagger}$) necessary for all steps. The detailed analyses leading to this result can be found in Appendix B.
@@ 3 Resources comparison
In Tab. 1, we compare the required resources for the QEE and CPS methods: the required coherence time and the number of state preparation circuits. The coherence time is quantified by the duration of the state preparation circuit, $t_{prep}$, which can also be seen as a measure for the maximum circuit depth. The scaling of the number of state preparations is the proper indicator for the complexity of the CPS method when the cost of implementing controlled reflection and Pauli operations does not scale faster than that of state preparation. This requirement is satisfied easily for the controlled Pauli operation when the $\hat{P}_{j}$ are $k$-local. On the other hand, the controlled reflection operation can be implemented using resources that scales linearly in the number of qubits on the Rydberg atom platform [25, 12, 43, 38].
@@
The QEE only requires a coherence time of the multi-qubit device of $\sim t_{prep}$ since after each state preparation, the qubits are measured. In comparison, the CPS requires a modest increase in the coherence time of the quantum device that scales logarithmic with the number of Pauli strings and the target variance, $\eta$. In return, the CPS provides a much better estimate of $\langle O\rangle$ than the QEE for a fixed number of state preparations. We find that
$\frac{\sigma^{2}{(\hat{O}_{CPS})}}{\sigma^{2}{(\hat{O}_{QEE})}}\sim\frac{1}{N},$ (4)
@@
which shows that the CPS achieves a Heisenberg-like scaling of the variance in the number of Pauli strings compared to the standard quantum limit scaling of the QEE.
@@
The limiting factor of both QEE and QSP will be the accumulation of operational errors in the final estimate of the observable. For both approaches, gate errors will reduce the accuracy of the final estimate of the observable. This reduction will have the same linear dependence on the number of Pauli strings in the observable decomposition for both approaches (see Appendix C). For the estimation of observable such as the energy of a $\text{H}_{3}^{+}$ molecule, which can be decomposed into $N\sim 59$ Pauli strings [5] on $4$ qubits, an estimate of the energy with a variance of $\eta\sim 1$ could be obtained for gate error probabilities around $\sim 0.5\cdot 10^{-3}$ and $10^{2}$ state preparations using CPS. In comparison, QEE would require $\sim 10^{3}$ state preparations to reach a similar variance. We do note, however, that the QEE approach would only require gate error probabilities on the order of $5\cdot 10^{-3}$ (we refer to Appendix D for more details). To simulate larger molecules, the performance of current hardware needs to be improved for both our method and QEE. For example, the CPS simulation of a LiH molecule with a decomposition of $\sim 630$ Pauli strings on $10$ qubits, requires $\sim 10^{3}$ state preparations and gate error probabilities $\sim 10^{-5}$, while the QEE method requires $\sim 10^{5}$ state preparations and gate error probabilities $\sim 10^{-4}$ to achieve the same variance of $\eta\sim 1$. Note, however, that the estimates for the required gate error probabilities are conservative since we assuming that just a single error corrupts the estimation. For specific implementations and hardware models better bounds can likely be obtained.

### v4

@@ Abstract
The estimation of many-qubit observables is an essential task of quantum information processing. The generally applicable approach is to decompose the observables into weighted sums of multi-qubit Pauli strings, i.e., tensor products of single-qubit Pauli matrices, which can readily be measured with low-depth Clifford circuits. The accumulation of shot noise in this approach, however, severely limits the achievable variance for a finite number of measurements. We introduce a novel method, dubbed coherent Pauli summation (CPS), that circumvents this limitation by exploiting access to a single-qubit quantum memory in which measurement information can be stored and accumulated. CPS offers a reduction in the required number of measurements for a given variance that scales linearly with the number of Pauli strings in the decomposed observable. Our work demonstrates how a single long-coherence qubit memory can assist the operation of many-qubit quantum devices in a cardinal task.
@@ 1 Introduction
One approach for observable estimation with minimal sampling complexity is the quantum phase estimation (QPE) algorithm [19, 18, 40, 29, 45, 32]. Its implementation, however, requires qubit systems with low noise and long coherence times for high precision estimation since the measurement circuit depth is inversely proportional to the square root of the achievable estimation variance. In addition, the observable of interest has to be encoded as a unitary transformation which is, in general, a non-trivial requirement.
As an alternative, the quantum energy (expectation) estimation (QEE) approach of decomposing the observable into a weighted sum of $N$ multi-qubit Pauli strings is commonly used in the variational quantum eigensolver [34]. While QEE minimizes the measurement circuit depth by requiring only a single layer of single-qubit rotations, it also suffers from increased sample complexity due to the accumulation of shot noise in the estimation procedure. Specifically, the expectation values of Pauli strings are estimated independently, and the observable is then calculated as a linear combination of these. Consequently, to estimate an observable comprised of $N$ Pauli stings to a variance $\eta$, each Pauli string should be estimated to a variance $O(\eta/N)$ resulting in an overall sample complexity scaling as $O(N^{2})$. This accumulation of noise poses a “shot noise bottleneck" since the amount of measurements will ultimately be limited by the available runtime of the device before, e.g., re-calibration of the device is needed. The measurement process itself is also often one of the most time consuming operations in current quantum devices [17, 5, 38].
@@
Here, we propose a novel approach dubbed coherent Pauli summation (CPS) that overcomes the shot-noise bottleneck through the use of a single-qubit quantum memory (QM) and Quantum Signal Processing (QSP) techniques [24, 10]. Our method allows for a direct measurement of the multi-qubit observable by estimating the phase of the memory qubit at the end of the protocol. Hence, the accumulation of shot noise, originating from the summation of individually estimated mean values of Pauli strings in the QEE approach, is prevented. Importantly, the properties of the QSP-based encoding allow this task to be completed with processing qubits whose coherence time scales only logarithmically with the required estimation variance, in contrast to the linear scaling of the QPE algorithm.
The performance of the QPE algorithm for observable estimation can significantly vary depending on the method. For example, in Refs. [8, 13, 18, 45], Heisenberg limited scaling is achieved assuming the observable can be encoded with a unitary evolution. This assumption is difficult to satisfy in practice and limits the applicability of the result especially for current stage of technological development. Other types of QPE algorithms relax this restrictive assumption but result in worse performance [32, 43].
The critical point of the CPS method is to employ QSP techniques to encode the mean value of Pauli strings in the phase of a single qubit’s quantum memory, enabling us to artificially rescale the phase between 00 and $2\pi$ without additional cost [24]. Furthermore, this allows us to obtain Heisenberg scaling of the estimation variance with circuit depths on the many-qubit device that increase only logarithmically with the inverse of the achievable estimation variance. This is because the many-qubit device can be readout and reset after each encoding into the memory qubit. The total circuit depth of the memory qubit is inversely proportional to the square root of the achievable estimation variance as in standard QPE. From this perspective, one can think of CPS as a new variation of QPE, using QSP and the single qubit QM to achieve the Heisenberg scaling without paying a high price of a unitary preparation or long circuits depths on the many-qubit device. Thus, CPS outlines how a single error corrected qubit memory can advance the performance of larger scale quantum devices. The lack of shot noise accumulation results in a gain in the variance of the estimate of $O(1/\sqrt{N})$ compared to performing QPE of individual Pauli strings which are summed classically [43] and the QPE scaling results in a gain of $O(1/T)$ where $T$ is the amount of state preparations compared to QEE.
@@ 2 CPS method
In general, the state $|\Psi\rangle$ is repeatedly prepared, and the Pauli strings are measured sequentially via projective measurements to obtain estimates of every Pauli string: $\langle\hat{P}_{j}\rangle\equiv\langle\Psi|\hat{P}_{j}|\Psi\rangle$ [34, 27]. The mean value of the observable $\hat{O}$ can be calculated classically after estimating all $\langle\hat{P}_{j}\rangle$, $j=[1,N]$. This approach, however, suffers from the accumulation of shot noise from the individually estimated mean values of the Pauli strings, as described above. If each estimate $\langle\hat{P}_{j}\rangle$ is estimated with a variance $\sigma^{2}(\langle\hat{P}_{j}\rangle)=\eta_{p}$, the variance of the final estimate is $\sigma^{2}(\hat{\langle O\rangle})\approx N\eta_{p}$, assuming roughly equal weights of the $\hat{P}_{j}$’s in (1).
@@
    Obtain rough estimates of the mean values of every Pauli string $\langle\Psi_{0}|\hat{P}_{j}|\Psi_{0}\rangle=\langle\hat{P}_{j}\rangle$, $j=[1,N]$ by performing $O\left(\log{\left(\frac{N}{\sqrt{\eta}}\right)}\right)$ projective measurements similar to the QEE approach. This step estimates $s_{j}=\text{sign}{(a_{j}\langle\hat{P}_{j}\rangle)}$ i.e. the sign information of the Pauli strings.
@@
    In the second step, $\tau_{j}|\langle\hat{P}_{j}\rangle|$, $\tau_{j}\equiv s_{j}\epsilon|a_{j}|$ is directly encoded in the phase of the single qubit memory using a modified phase-kickback algorithm [21] together with QSP techniques. Here, $\epsilon$ is a tuning parameter that is used to circumvent the general "modulo $2\pi$" ambiguity of phase estimation, without a substantial scaling of the QSP circuit depth, which we will detail below. The encoding is done sequentially for all Pauli strings in the decomposition, resulting in a final phase of $\sim\langle\hat{O}\rangle$ followed by a projective measurement of the single qubit state.
@@
    The previous step is repeated a number of times with a varying parameter $\epsilon$ to obtain the final estimate of $\langle\hat{O}\rangle$.
At the end of these three steps, $\langle{O}\rangle$ can be estimated up to a variance of $\eta$ with an overall sample complexity of $T=O\left(N/\sqrt{\eta}(\frac{\log{(N/\sqrt{\eta})}}{\log{\log{(N/\sqrt{\eta})}}})\right)$. We have counted this scaling as the number of state preparation circuits $V$ required in the method, which also includes the extra state preparation circuits that are part of the QSP step as detailed below.
@@
Figure 2: Quantum circuit realizing the QSP to encode one Pauli by a sequence of controlled unitaries $\hat{U}_{\phi_{i}}(\theta_{j})$, $i=[1,n]$, built from Hadamard gates, $\hat{R}_{\phi}=e^{\frac{i\phi\hat{\sigma}_{z}}{2}}$, and controlled $\hat{U}_{P_{j}}$. |$\ket{\psi}_{m}=\alpha\ket{+}_{m}+\beta\ket{-}_{m}$ denotes the state of a single memory qubit. The state of the mutli-qubit device is reset after each encoding, resulting in modest requirements on the coherence time of the qubits.
Method Number of state preparations Qubits Coherence time
QEE $O(N^{2}/\eta)$ Processing qubits $t_{prep}$
QPEh $O({N}^{3/2}/\sqrt{\eta})$ Processing qubits $t_{prep}O(\sqrt{N}/\sqrt{\eta}))$
CPS $O\left(\frac{N}{\sqrt{\eta}}\frac{\log{(N/\sqrt{\eta})}}{\log{\log{(N/\sqrt{\eta})}}}\right)$ Processing qubits $t_{prep}O\left(\frac{\log{(N/\sqrt{\eta})}}{\log{\log{(N/\sqrt{\eta})}}}\right)$
Memory qubit $t_{prep}O\left(\frac{N}{\sqrt{\eta}}\frac{\log{(N/\sqrt{\eta})}}{\log{\log{(N/\sqrt{\eta})}}}\right)$
Table 1: Comparison of resources for the QEE, QPE with the Heisenberg scaling per Pauli and CPS methods for a fixed target variance $\eta$. We have quantified the necessary coherent time of the processing qubits of quantum device and the memory qubit in terms of the time required for a single state preparation $t_{prep}$, which is assumed to be the dominant scale. The observable is assumed to be decomposed into a summation of $N$ Pauli strings (1). We refer to the main text and Appendix C for further details)
@@ 2.1 Phase wrapping
Measuring the phase in Eq. (3) by a series of projective measurements would result in an estimation variance scaling as the standard quantum limit instead of the Heisenberg limit (see Appendix A). Moreover, since the phase encoding only provides an estimate of $\langle\hat{O}\rangle$ mod $2\pi$, we are facing a phase wrapping problem. The observable we want to estimate can be written as $2\pi C+B$, where $C\in\mathbb{Z}$, $B\in\mathbb{R}$. So direct measurement will provide us with a precise estimate of $B$, losing the information about the amount of phase wraps $C$. To mitigate this issue and to achieve the Heisenberg scaling, we have included the factor of $\epsilon$ in the encoding. This brings us to the third and final step of the CPS. To control the phase wrapping, we use the sampling approach introduced in Refs. [13, 18] to estimate $C$ and $B$ separately. Instead of using fixed $\epsilon$ to encode $\theta_{j}$, we sample at multiple orders ${\epsilon}_{l}(\epsilon_{0})\equiv 2^{l-1}\epsilon_{0}$, $l=1,2,\dots$ to gradually enclose on $\langle\hat{O}\rangle$. This trick can be seen as sequentially estimating the digits of $\langle\hat{O}\rangle$.
For estimating $C$, the parameter $\epsilon_{0}\sim 1/N$ is chosen such that the condition $\sum_{j=1}^{N}\tau_{j}\cos{\theta_{j}}\leq 2\pi$ is satisfied. The QSP procedure with different ${\epsilon}_{l}(\epsilon_{0})$ is repeated for $l\in[1,L]$, where $L=O(\log(1/\sqrt{\eta}))$. The estimation of $B$ is performed in a similar way, but with $\epsilon_{0}=1$. Note that naively scaling the variable down by a factor of $\epsilon_{0}\sim 1/N$ and conducting standard phase estimation would require number of state preparations that scale with $N^{2}$. However, because we encode each Pauli string as a phase $e^{i\theta}$ with $\theta\in\{0,2\pi\}$ and QSP techniques allow us to realize maps between $\{0,2\pi\}\rightarrow\{0,2\pi\}$ with a nearly constant number of QSP iterations, we can estimate the observable $\hat{O}$ up to $O(1)$ variance using only $\log(N)$ state preparations for each Pauli string. Consequently, we find that with a total number of state preparations scaling as $O\left(\frac{N}{\sqrt{\eta}}\frac{\log{(N/\sqrt{\eta})}}{\log{\log{(N/\sqrt{\eta})}}}\right)$, the CPS method results in an estimate of $\hat{O}$ with a variance of $\eta$.
@@ 3 Resources comparison
In Tab. 1, we compare the required resources for the QEE, QPEh, and CPS methods: the required coherence time of the processing qubits and the number of state preparation circuits. Here QPEh referes to estimating each Pauli string with QPE and then sum the estimates classically to obtain $\langle O\rangle$ [8, 18, 43]. The coherence time is quantified by the time required for the state preparation circuit, $t_{prep}$. The scaling of the number of state preparations is the proper indicator for the complexity of the CPS method when the cost of implementing controlled reflection and Pauli operations does not scale faster than that of state preparation. This requirement is satisfied easily for the controlled Pauli operation when the $\hat{P}_{j}$ are $k$-local. On the other hand, the controlled reflection operation can be implemented using resources that scales linearly in the number of qubits on the hardware that supports native multi-qubit gates [28, 15, 47, 41].
@@
The QEE method only requires a coherence time of the multi-qubit device of $\sim t_{prep}$ since after each state preparation the qubits are measured. The QPEh method requires a coherence time scaling as $t_{prep}O(\sqrt{N}/\sqrt{\eta}))$ since each Pauli string is estimated using standard QPE. In comparison, the CPS method comes with different requirements on coherence time for the multi-qubit device and the single qubit memory. For the multi-qubit device, a modest increase in the coherence time compared to QEE which scales as $O\left(\frac{\log{(N/\sqrt{\eta})}}{\log{\log{(N/\sqrt{\eta})}}}\right)$ is required. This comes from the requirement of encoding each Pauli string using a logarithmically scaling number of QSP steps. On the other hand, the coherence time of the memory qubit has a similar scaling as QPEh up to a logarithmic factor. In return, the CPS provides a much better estimate of $\langle O\rangle$ than the QEE for a fixed number of state preparations. We find that
$\frac{\sigma^{2}{(\hat{O}_{CPS})}}{\sigma^{2}{(\hat{O}_{QEE})}}\sim\frac{1}{T},$ (4)
@@
which shows that the CPS achieves a Heisenberg-like scaling of the variance in the number of state preparations compared to the standard quantum limit scaling of the QEE. CPS outperforms the QPEh too:
$\frac{\sigma^{2}{(\hat{O}_{CPS})}}{\sigma^{2}{(\hat{O}_{QPEh})}}\sim\frac{1}{N},$ (5)
which is because CPS circumvents the accumulation of noise in the classical summation in QPEh.
@@
The limiting factor of both QEE and CPS will be the accumulation of operational errors in the final estimate of the observable. For both approaches, gate errors will reduce the accuracy of the final estimate. This reduction will have the same linear dependence on the number of Pauli strings in the observable decomposition for both approaches (see Appendix D). For the estimation of observable such as the energy of a $\text{H}_{3}^{+}$ molecule, which can be decomposed into $N=59$ Pauli strings [6] on $4$ qubits, an estimate of the energy with a variance of $\eta=1$ could be obtained for gate error probabilities $\sim 4.5\cdot 10^{-5}$ and at least $1.1\times 10^{4}$ state preparations using CPS. In comparison, QEE would require $\sim 3.48\times 10^{3}$ state preparations and QPEh $\sim 4.85\times 10^{3}$ to reach a similar variance. We do note, however, that the QEE approach would only require gate error probabilities on the order of $4.2\times 10^{-3}$, while QPEh $\ 1.8\times 10^{-4}$ (we refer to Appendix E for more details). To simulate larger molecules, the performance of current hardware needs to be improved for both our method and QEE. For example, the CPS simulation of a LiH molecule with a decomposition of $\sim 630$ Pauli strings on $10$ qubits, requires $\sim 1.4\times 10^{5}$ state preparations and gate error probabilities $\sim 1.43\times 10^{-6}$, while the QEE method requires $\sim 3.96\times 10^{5}$ state preparations and gate error probabilities $\sim 1.6\times 10^{-4}$ to achieve the same variance of $\eta=1$. The QPEh requires $1.69\times 10^{4}$ state preparations and gate error probabilities of a rate $\sim 2.1\times 10^{-6}$. Note, however, that the estimates for the required gate error probabilities are conservative since we assume that just a single error corrupts the estimation. For specific implementations and hardware models better bounds can likely be obtained.